For the hydrolysis of a salt of a weak acid and a weak base, show that
K h = K w/K a K b
1
Expert's answer
2020-02-25T04:46:29-0500

hydrolysis of NH4HS salt (ammonium hydrosulfide) - a salt of a weak acid and a weak base


NH4+ + HS + НОН ⇆ NH4ОН + H2S

Ка(H2S)=1.02*10-7

Кb(NH4ОН )=1.76*10-5

For a salt of a weak monoacid base and a weak monobasic acid:

Кг=Кw / (Ка ∙ Кb) = 10-14 / (1.02*10-7 *1.76*10-5) = 5.570*10-3